OpenStudy (anonymous):

can someone help? The length of a rectangle is equal to 3 more than 3 times the width. The perimeter is 54 ft. What are the length and width of the rectangle?

OpenStudy (igreen):

2 times the length plus 2 times the width equals perimeter therefore, 2L+2W= P the length is three times the width therefore, L = 3W and you know the perimeter is 96 and inserting 3w for length you get, 2(3W)+2W=96, then 6W+2W=96, then 8W=96. Divide both sides by 8 and you get W=12. The width is 12. And the length is 36. http://www.answers.com/Q/The_length_of_a_rectangle_is_3_times_the_width_the_perimeter_is_96cm_Find_the_width

OpenStudy (anonymous):

thx